# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v1

@@ 4.4 HRM Generalization Across Different Domains
To broaden the applicability of HRM and evaluate its generalization capability, we assess HRM and PRM, trained on the PRM800K dataset, on the Math500 [7] and GSM8K [6] datasets. Table 6 shows that HRM demonstrates superior generalization in Math500 compared to PRM, indicating its ability to handle complex mathematical reasoning tasks.
However, the performance difference between HRM and PRM in GSM8K is marginal, as the GSM8K dataset primarily consists of relatively simple arithmetic problems. A strong policy model can typically solve these problems within three steps on the GSM8K dataset, limiting the impact of HRM’s advantages, such as its ability to assess multi-step reasoning coherence and facilitate self-reflection through its training data. Nevertheless, as shown in Table 6, HRM still slightly outperforms PRM on simpler datasets like GSM8K. More importantly, in more complex datasets such as Math500, HRM demonstrates significantly greater robustness compared to PRM, highlighting its effectiveness in handling challenging reasoning tasks.
@@ 5 Conclusion
In this paper, we present Hierarchical Reward Model, which enhances multi-step reasoning evaluation by integrating fine-grained and coarse-grained assessments, improving reasoning coherence and self-reflection. We further introduce Hierarchical Node Compression, a data augmentation method that optimizes MCTS-based autonomous annotation, enhancing label diversity while expanding training data with minimal computational cost. Extensive experiments on PRM800K demonstrate HRM’s superior robustness over PRM, with strong generalization across GSM8K and MATH500. Additionally, MCTS-generated auto-labeled data enables policy model fine-tuning, further improving reasoning performance.
@@ A.3 Self-Training
The logarithmic scaling and weighting factor $\lambda$ are added to balance the impact of KL divergence. Without these adjustments, KL loss would range from 0 to 8000, while the language modeling loss remains between 0 and 12, leading to an imbalance. The logarithm ensures a more stable contribution of KL loss during training.

### v2

@@ 4.4 HRM Generalization Across Different Domains
To broaden the applicability of HRM and evaluate its generalization capability, we assess HRM and PRM, trained on the PRM800K dataset, on the Math500 [7] and GSM8K [6] datasets. Table 6 shows that HRM exhibits greater robustness across different domains, demonstrating superior generalization performance, particularly in Math500, where it effectively handles complex mathematical reasoning tasks.
However, the performance difference between HRM and PRM on the GSM8K dataset is marginal, as GSM8K primarily consists of relatively simple arithmetic problems. A strong policy model can typically solve these problems within three steps, reducing the impact of HRM’s key advantages, such as assessing multi-step reasoning coherence and facilitating self-reflection. Nevertheless, as shown in Table 6, HRM still achieves a slight performance edge over PRM, even on simpler datasets like GSM8K.
@@ 5 Conclusion
In this paper, we present Hierarchical Reward Model, which enhances multi-step reasoning evaluation by integrating fine-grained and coarse-grained assessments, improving reasoning coherence and self-reflection. We further introduce Hierarchical Node Compression, a data augmentation method that optimizes MCTS-based autonomous annotation, enhancing label diversity while expanding training data with minimal computational cost. Extensive experiments on PRM800K dataset demonstrate HRM’s superior robustness over PRM, with strong generalization across GSM8K and MATH500 dataset. Additionally, MCTS-generated auto-labeled data enables policy model fine-tuning, further improving reasoning performance.
@@ A.3 Self-Training
The logarithmic scaling and weighting factor $\lambda$ are added to balance the impact of KL divergence. Without these adjustments, KL loss would range from 0 to 20000, while the language modeling loss remains between 0 and 12, leading to an imbalance. The logarithm ensures a more stable contribution of KL loss during training.